OpenStudy (anonymous):

what is the perpendicular to -7x/6 that passes through point (-8,13)

OpenStudy (anonymous):

the opposite inverse of the slope is the slope of your perpendicular rect, so y=6x/7+q, sobstitute x and y => 13=-48/7+q => q=13+48/7 soooo y=6x/7+ (13+48/7)

OpenStudy (ricardovg):

In two lines, perpendicular means m1*m2=-1 -(7/6)*(6/7)=-1 then m2=(6/7) (y-13)/(x+8)=6/7 => y=(6/7)x+139/7

OpenStudy (anonymous):

@ ricardovg...how did you get 139?

OpenStudy (ricardovg):

(y-13)/(x+8)=6/7 (y-13)=(6/7)(x+8) y - 13= (6/7)x+(6/7)*8 y = (6/7)x+(48/7) + 13 y = (6/7)x + (13*7+48)/7 = (6/7)x + (139)/7 ,here it is 139